# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ 4.1. The main theorems
We now remark that the space $\langle\omega_{1},\dots,\omega_{n-1}\rangle_{{\mathbb{F}_{p}}}$ is a space $L_{\lambda p^{n-1},n-1}$ and then by inductive hypothesis there exist $S_{1},\dots,S_{n-1}\in k[X]$ of degree $\lambda p$ such that $\frac{P_{i}}{P}=(-1)^{i-1}\frac{\Delta_{n-2}(\underline{\hat{S_{i}}})}{\Delta_{n-1}(\underline{S})}$, which means that $P_{i}=(-1)^{i-1}\Delta_{n-2}(\underline{\hat{S_{i}}})\frac{P}{\Delta_{n-1}(\underline{S})}$. We then have that
@@
To conclude, we need to show that there exist $Q_{1},Q_{2},\cdots,Q_{n}\in k[X]$ of degree $\lambda$ such that $P=\Delta_{n}(\underline{Q})$ and $P_{i}=(-1)^{i-1}\Delta_{n-1}(\underline{\hat{Q}_{i}})$. Let $\varphi:k^{n}\longrightarrow k^{n}$ be the map defined by $\left(\varphi(\underline{a})\right)_{i}=(-1)^{i-1}\Delta_{n-1}(\underline{\hat{a_{i}}})$. Then by Proposition A.15 (since $\Delta_{n}(\underline{P})\neq 0$) there exist $n$-tuples $\underline{Q},\underline{R}$ of elements of $k(X)^{\mathrm{alg}}$ satisfying $\varphi(\underline{R})=\underline{Q}$, $\varphi(\underline{Q})=\underline{P}$, and
@@
which is a polynomial of degree $\lambda$ thanks to the fact that the zeroes of $\Delta_{n-1}(\underline{S})$ are simple and correspond to the set of poles of the space $\langle\omega_{1},\dots,\omega_{n-1}\rangle_{{\mathbb{F}_{p}}}$ (see Theorem 4.7). Moreover, by Lemma 2.4, we have that $\deg(Q_{n})=\lambda$. In a completely analogous way, we can show that the $Q_{i}$’s are polynomials of degree $\lambda$ also for $11\leq i\leq n-$ 1. ∎
@@
Let $\Omega=\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ be a space $L_{\lambda p^{n-1},n}$ and let $Q_{1},\dots,Q_{n}$ be the $n$-uple of polynomials arising from Theorem 4.8. Then, for every $1\leq t\leq n$ the subspace $\Omega_{t}=\langle\omega_{n-t+1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ is such that
@@
Let $\Omega=\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ be a space $L_{\lambda p^{n-1},n}$ and let $Q_{1},\dots,Q_{n}$ be the $n$-uple of polynomials arising from Theorem 4.8 for this basis.
Denote by $q_{i}$ the leading coefficient of $Q_{i}$. Then we have the equality of Moore determinants
@@
$\alpha=\frac{\Delta_{n}(q_{1},\dots,q_{n})}{\Delta_{t}(P_{\underline{q}}(q_{n-t+1}),\dots,P_{\underline{q}}(q_{n}))\Delta_{n-t}(q_{1},\dots,q_{n-t})}\in k^{\times}$
and $P_{\underline{q}_{t}}$ the structural polynomial of the vector space $\langle q_{1},\dots,q_{n-t}\rangle_{{\mathbb{F}_{p}}}$.
@@
We know that the zeroes of $\Delta_{n}(\underline{Q})$ are simple and consist of the set ${\mathcal{P}}(\Omega)$, and by Proposition 4.6 the zeroes of $\Delta_{t}(P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,P_{{\mathcal{Q}}_{n-t}}(Q_{n}))$ are simple and consist of the set ${\mathcal{P}}(\Omega_{t})$ with $\Omega_{t}=\leavevmode\nobreak\ \langle\omega_{n-t+1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$. We may then apply Corollary 4.15 to see that the set of zeroes of the polynomials on both sides of the equation are equal, and that these zeroes are all simple. The corollary then follows from a comparison of the leading coefficients of these polynomials. ∎
We conclude this part with a result that describes the $n$-uples of polynomials giving rise to a basis of an étale pullback of a space $L_{\lambda p^{n-1},n}$.
@@
[…]
@@
4.2. Polynomial conditions for the existence of spaces $L_{\lambda p^{n-1},n}$
The results of Theorems 4.7 and 4.8 show that the existence of a space $L_{\lambda p^{n-1},n}$ is equivalent to the existence of a $n$-uple $\underline{Q}\in k[X]^{n}$ satisfying certain conditions. In analogy with the case $n=2$, Proposition 2.14 (iii) gives us a way to check this by solving the following system of polynomial equations (note that in this situation the number of poles is $\lambda p^{n-1}$, and the number of equations is computed accordingly)

### v2

@@ 4.1. Constructing spaces $L_{\lambda p^{n-1},n}$ from polynomials
We now remark that the space $\langle\omega_{1},\dots,\omega_{n-1}\rangle_{{\mathbb{F}_{p}}}$ is a space $L_{\lambda p^{n-1},n-1}$ and then by inductive hypothesis there exist $S_{1},\dots,S_{n-1}\in k[X]$ of degree $\lambda p$ such that $\frac{P_{i}}{P}=(-1)^{i-1}\frac{\Delta_{n-2}(\underline{\hat{S_{i}}})}{\Delta_{n-1}(\underline{S})}$, which means that $P_{i}=(-1)^{i-1}\Delta_{n-2}(\underline{\hat{S_{i}}})\frac{P}{\Delta_{n-1}(\underline{S})}$ for every $1\leq i\leq n-1$. We then have that
@@
[…]
@@
which is a polynomial of degree $\lambda$ thanks to the fact that the zeroes of $\Delta_{n-1}(\underline{S})$ are simple and correspond to the set of poles of the space $\langle\omega_{1},\dots,\omega_{n-1}\rangle_{{\mathbb{F}_{p}}}$ (see Theorem 4.7). Moreover, by Lemma 2.4, we have that $\deg(Q_{n})=\lambda$. In a completely analogous way, we can show that the $Q_{i}$’s are polynomials of degree $\lambda$ also for $1\leq i\leq n-1$. Finally, we need to multiply the $Q_{i}^{\prime}s$ by a constant in order for them to satisfy the Theorem. In fact, we have
$\Delta_{n}(\underline{P})=\Delta_{n}(\Delta_{n-1}(\underline{\hat{Q}_{1}}),\dots,(-1)^{i-1}\Delta_{n-1}(\underline{\hat{Q}_{i}}),\dots,(-1)^{n-1}\Delta_{n-1}(\underline{\hat{Q}_{n}})),$
which combined with (4.11) and Theorem A.3 gives
$P^{1+p+\cdots+p^{n-2}}=\Delta_{n}(\underline{Q})^{1+p+\cdots+p^{n-2}}.$
Then, there exists $\theta^{\prime}\in k$ with $\theta^{\prime 1+p+\cdots+p^{n-2}}=1$ and $\Delta_{n}(\underline{Q})=\theta^{\prime}P$.
Let $\theta^{\prime\prime}\in k$ with $(\theta^{\prime\prime})^{p^{n-1}}\theta^{\prime}=1$. This satisfies $(\theta^{\prime\prime})^{1+p+\cdots+p^{n-2}}=1$ and then by Proposition A.16 we have $\varphi(\theta^{\prime\prime}\underline{Q})=\underline{P}$ and
$\Delta_{n}(\theta^{\prime\prime}\underline{Q})=(\theta^{\prime\prime})^{1+p+\cdots+p^{n-1}}\Delta_{n}(\underline{Q})=(\theta^{\prime\prime})^{p^{n-1}}\theta^{\prime}P=P.$
@@
Let $\Omega=\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ be a space $L_{\lambda p^{n-1},n}$ and let $Q_{1},\dots,Q_{n}$ be the $n$-tuple of polynomials arising from Theorem 4.10. Then, for every $1\leq t\leq n$ the subspace $\Omega_{t}=\langle\omega_{n-t+1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ is such that
@@
Let $\Omega=\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ be a space $L_{\lambda p^{n-1},n}$ and let $Q_{1},\dots,Q_{n}$ be the $n$-tuple of polynomials arising from Theorem 4.10 for this basis.
Denote by $q_{i}$ the leading coefficient of $Q_{i}$. Then, for every $1\leq t\leq n$, we have the equality of Moore determinants
@@
$\alpha=\frac{\Delta_{n}(q_{1},\dots,q_{n})}{\Delta_{t}(P_{\underline{q}}(q_{n-t+1}),\dots,P_{\underline{q}}(q_{n}))\Delta_{n-t}(q_{1},\dots,q_{n-t})}\in k^{\times},$
and $P_{\underline{q}_{t}}$ is the structural polynomial of the vector space $\langle q_{1},\dots,q_{n-t}\rangle_{{\mathbb{F}_{p}}}$.
@@
We know that the zeroes of $\Delta_{n}(\underline{Q})$ are simple and consist of the set ${\mathcal{P}}(\Omega)$, and by Proposition 4.6 the zeroes of $\Delta_{t}(P_{{\mathcal{Q}}_{n-t}}(Q_{n-t+1}),\dots,P_{{\mathcal{Q}}_{n-t}}(Q_{n}))$ are simple and consist of the set ${\mathcal{P}}(\Omega_{t})$ with $\Omega_{t}=\penalty 10000\ \langle\omega_{n-t+1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$. We may then apply Corollary 4.17 to see that the set of zeroes of the polynomials on both sides of the equation are equal, and that these zeroes are all simple. The corollary then follows from a comparison of the leading coefficients of these polynomials. ∎
We conclude this part with a result that describes the $n$-tuples of polynomials giving rise to a basis of an étale pullback of a space $L_{\lambda p^{n-1},n}$.
@@
Let $\Omega=\langle\omega_{1},\dots,\omega_{n}\rangle_{{\mathbb{F}_{p}}}$ be a space $L_{\lambda p^{n-1},n}$ and let $Q_{1},\dots,Q_{n}$ be the $n$-tuple of polynomials arising from Theorem 4.10. Let $S(X)\in k[X]$ with $S^{\prime}(X)\in k^{\times}$ and let $\sigma^{\star}(\Omega)$ be the pullback of $\Omega$ with respect to the morphism ${\mathbb{P}}^{1}_{k}\overset{\sigma}{\rightarrow}{\mathbb{P}}^{1}_{k}$ induced by $X\mapsto S(X)$ (cf. Lemma 2.14). Then the polynomials arising from Theorem 4.10 for $\sigma^{\star}(\Omega)$ are $\big(\eta Q_{1}(S),\dots,\eta Q_{n}(S)\big)$ with $\eta^{p^{n-1}}=\frac{1}{S^{\prime}(X)}$. In particular, we have that ${\mathcal{P}}(\sigma^{\star}(\Omega))=\{a\in k\ |\ S(a)\in{\mathcal{P}}(\Omega)\}$.
@@
4.2. Dickson invariants and spaces $L_{\lambda p^{n-1},n}$
In the previous section, we proved several results using the structural polynomial of the vector space generated by the $Q_{i}$’s. Dickson invariants (see Appendix A) provide an elegant framework for describing this polynomial and can be effectively applied to the problem of determining the existence of spaces $L_{\lambda p^{n-1},n}$. First of all, Dickson invariants can be used to give a shorter and potentially more insightful proof of Theorem 4.7. To show this, let us adopt the same notation as in that theorem and recall that the non-trivial part is to show that if the form $\omega_{n}$ is logarithmic, then the forms $\omega_{i}$ are logarithmic for every $i=1,\dots,n$. Equivalently, we need to prove that if $\left(\frac{P}{P_{n}}\right)^{p-1}$ satisfies the system of equations (4.9), then so does $\left(\frac{P}{P_{i}}\right)^{p-1}$ for every $i$.
From (A.6(i)), it follows that
$\left(\frac{P}{P_{n}}\right)^{p-1}=\left(\frac{\Delta_{n}(\underline{Q})}{\Delta_{n-1}(\underline{\hat{Q}_{n}})}\right)^{p-1}=c_{n,n-1}-c_{n-1,n-2}^{p}\in{\mathbb{F}}_{p}[Q_{1},\dots,Q_{n}].$
We then consider the element $\tau_{(n,i)}\in\operatorname{GL}_{n}({\mathbb{F}}_{p})$, acting on ${\mathbb{F}}_{p}[Q_{1},\dots,Q_{n}]$ by exchanging $Q_{n}$ with $Q_{i}$ and leaving all other variables unchanged. Applying the action of $\tau_{(n,i)}$ to $\left(\frac{P}{P_{n}}\right)^{p-1}$, we obtain:
$\left(\frac{\Delta_{n}(\underline{Q})}{\Delta_{n-1}(\underline{\hat{Q}_{i}})}\right)^{p-1}=\left(\frac{\tau_{(n,i)}\star\Delta_{n}(\underline{Q})}{\tau_{(n,i)}\star\Delta_{n-1}(\underline{\hat{Q}_{n}})}\right)^{p-1}=-(\tau_{(n,i)}\star c_{n,n-1})+\left(\tau_{(n,i)}\star c_{n-1,n-2}\right)^{p}.$
Since $c_{n,n-1}$ is a Dickson invariant, we have $\tau_{(n,i)}\star c_{n,n-1}=c_{n,n-1}$, and it follows that
$\left(\frac{P}{P_{i}}\right)^{p-1}-\left(\frac{P}{P_{n}}\right)^{p-1}\in\left({\mathbb{F}}_{p}[Q_{1},\dots,Q_{n}]\right)^{p}\subset\left({\mathbb{F}}_{p}[X]\right)^{p}.$
In particular, $\left(\frac{P}{P_{n}}\right)^{p-1}$ satisfies (4.9) if and only if $\left(\frac{P}{P_{i}}\right)^{p-1}$ does.
More crucially, the identities in Proposition A.6 allow us to achieve two new key results: (1) a construction of a large class of previously unknown spaces $L_{\lambda 2^{n-1},n}$ in characteristic 2 and (2) a proof that spaces $L_{p^{n-1},n}$ (i.e., with $\lambda=1$) do not exist when $p>2$. The key step required for these two applications is a proposition that distills Theorems 4.7 and 4.10 into a new necessary and sufficient condition for the existence of a space $L_{\lambda p^{n-1},n}$ using the property of Dickson invariants.
Proposition 4.20.
Let $(Q_{1},\cdots,Q_{n})\in k[X]^{n}$ be a $n$-tuple of polynomials of degree $\lambda\geq 1$ with leading coefficients $q_{i}$ satisfying $\Delta_{n}(\underline{q})\neq 0$. Then, the $n$-tuple $\underline{Q}$ gives rise to a basis of a space $L_{\lambda p^{n-1},n}$ if and only if
$\left(\det(\underline{Q}^{\prime},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}})\cdot\det(\underline{Q},\underline{Q^{p}},\cdots,\underline{Q}^{p^{n-2}},\underline{Q}^{p^{n-1}})^{p-2}\right)^{(p-2)}=1,$
where the exponent (p2)(p-2) denotes the p2p-2-th derivative of a polynomial in $k[X]$.
Proof.
By A.6(ii), we have that $c_{n,n-1}=\frac{\widetilde{\Delta}_{n}(\underline{Q})}{\Delta_{n}(\underline{Q})}$, where
$\widetilde{\Delta}_{n}(\underline{Q}):=\det(\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}},\underline{Q}^{p^{n}}).$
Using this, we can compute the derivative $c_{n,n-1}^{\prime}$:
$c_{n,n-1}^{\prime}=\left(\frac{\widetilde{\Delta}_{n}(\underline{Q})}{\Delta_{n}(\underline{Q})}\right)^{\prime}=\frac{\widetilde{\Delta}_{n}(\underline{Q})^{\prime}\Delta_{n}(\underline{Q})-\widetilde{\Delta}_{n}(\underline{Q})\Delta_{n}(\underline{Q})^{\prime}}{\Delta_{n}(\underline{Q})^{2}}.$
Since $\Delta_{n}(\underline{Q})\neq 0$, the set of vectors $\{\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}},\underline{Q}^{p^{n-1}}\}$ is a basis of the $k(X)$-vector space $k(X)^{n}$ and therefore there exist $a_{0},\cdots a_{n-1},b_{0},\cdots,b_{n-1}\in k(X)$ such that
$\underline{Q}^{\prime}=\sum_{i=0}^{n-1}a_{i}\underline{Q}^{p^{i}}\;\;\;\mathrm{and}\;\;\;\underline{Q}^{p^{n}}=\sum_{i=0}^{n-1}b_{i}\underline{Q}^{p^{i}}.$
We first compute
$\Delta_{n}(\underline{Q})^{p}=\det(\underline{Q}^{p},\underline{Q}^{p^{2}},\cdots,\underline{Q}^{p^{n-1}},\underline{Q}^{p^{n}})=\det(\underline{Q}^{p},\underline{Q}^{p^{2}},\cdots,\underline{Q}^{p^{n-1}},\sum_{i=0}^{n-1}b_{i}\underline{Q}^{p^{i}})=(-1)^{n-1}b_{0}\Delta_{n}(\underline{Q}),$
and find that the coefficient $b_{0}$ is equal to $(-1)^{n-1}\Delta_{n}(\underline{Q})^{p-1}$. Then, by multilinearity of the determinant we find that
$\widetilde{\Delta}_{n}(\underline{Q})^{\prime}=\det(\underline{Q}^{\prime},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}},\underline{Q}^{p^{n}})=\det\left(\sum_{i=0}^{n-1}a_{i}\underline{Q}^{p^{i}},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}},\sum_{j=0}^{n-1}b_{j}\underline{Q}^{p^{j}}\right)=(a_{0}b_{n-1}-a_{n-1}b_{0})\Delta_{n}(\underline{Q}).$
With a similar computation, we get that
$\widetilde{\Delta}_{n}(\underline{Q})=\det\left(\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}},\sum_{j=0}^{n-1}b_{j}\underline{Q}^{p^{j}}\right)=b_{n-1}\Delta_{n}(\underline{Q})$
and
$\Delta_{n}(\underline{Q})^{\prime}=\det(\underline{Q}^{\prime},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}},\underline{Q}^{p^{n-1}})=\det\left(\sum_{i=0}^{n-1}a_{i}\underline{Q}^{p^{i}},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}},\underline{Q}^{p^{n-1}}\right)=a_{0}\Delta_{n}(\underline{Q}).$
By putting everything together, we conclude that $c_{n,n-1}^{\prime}=-a_{n-1}b_{0}=-a_{n-1}(-1)^{n-1}\Delta_{n}(\underline{Q})^{p-1}$. At the same time, we have
$\det(\underline{Q}^{\prime},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}})=\det\left(\sum_{i=0}^{n-1}a_{i}\underline{Q}^{p^{i}},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}}\right)=(-1)^{n-1}a_{n-1}\Delta_{n}(\underline{Q}).$
We then have
$c_{n,n-1}^{\prime}=-\det(\underline{Q}^{\prime},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}})\cdot\Delta_{n}(\underline{Q})^{p-2}.$
By Theorems 4.10 and 4.7 the $n$-tuple $\underline{Q}$ gives rise to a basis of a space $L_{\lambda p^{n-1},n}$ if and only if $(P_{n-1}(Q_{n})^{p-1})^{(p-1)}=-1$. By A.6(i), this is equivalent to $c_{n,n-1}^{(p-1)}=-1$ and the computation of the first derivative above allows us to rewrite the condition as
$\left[\det(\underline{Q}^{\prime},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{p^{n-2}})\cdot\Delta_{n}(\underline{Q})^{p-2}\right]^{(p-2)}=1,$
from which the result follows. ∎
4.2.1. Many new examples of spaces $L_{\lambda 2^{n-1},n}$
As a first application, we assume $p=2$ and give a construction for all values of $n\geq 2$ of new large classes of examples of spaces $L_{\lambda 2^{n-1},n}$ in characteristic $2$. In this case, Proposition 4.20 tells us that the necessary and sufficient condition for the existence of a space $L_{\lambda 2^{n-1},n}$ is the existence of a $n$-tuple of polynomials $\underline{Q}$ of same degree satisfying
$\det(\underline{Q}^{\prime},\underline{Q},\underline{Q}^{p},\cdots,\underline{Q}^{2^{n-2}})=1$
Remark 4.21.
Note that when $n=2$ this easily leads to a complete classification of spaces $L_{2\lambda,2}$. In fact, there are unique polynomials $U_{1},U_{2},V_{1},V_{2}\in k[X]$ such that $Q_{1}=U_{1}^{2}+XV_{1}^{2}$ and $Q_{2}=U_{2}^{2}+XV_{2}^{2}$. With these notations, we have $Q_{1}^{\prime}=V_{1}^{2}$ and $Q_{2}^{\prime}=V_{2}^{2}$, and hence
The condition imposed by Proposition 4.20 then is equivalent to
$V_{1}U_{2}+U_{1}V_{2}=1.$
If $\lambda$ is odd, then the polynomials $V_{1},V_{2}$ have degree $\frac{\lambda-1}{2}$. By Bézout’s theorem, if $V_{1}$ and $V_{2}$ are coprime, then there exists a unique pair $(U_{1},U_{2})$ with $\deg(U_{i})<\frac{\lambda-1}{2}$ satisfying the above condition. If we pick $(V_{1},V_{2})$ such that $V_{1}+V_{2}$ has also degree $\frac{\lambda-1}{2}$, and we let $(U_{1},U_{2})$ be the pair given by Bézout’s theorem, then by Proposition 4.20 the polynomials $U_{1}^{2}+XV_{1}^{2}$ and $U_{2}^{2}+XV_{2}^{2}$ give rise to a basis of a space $L_{2\lambda,2}$. If we relax the condition $\deg(U_{i})<\frac{\lambda-1}{2}$, then we have other pairs that satisfy Bezout’s theorem: these can be obtained from the minimal one $(U_{1},U_{2})$ as $(U_{1}+aV_{1},U_{2}+aV_{2})$ for $a\in k$.
If $\lambda$ is even, then the argument above works by choosing coprime $U_{1},U_{2}\in k[X]$ of degree $\frac{\lambda}{2}$ with $U_{1}+U_{2}$ of degree $\frac{\lambda}{2}$ and applying Bézout’s theorem to find $V_{1},V_{2}$. This shows that spaces $L_{2\lambda,2}$ exist for every $\lambda$ and in large abundance. While this is already known by [Pagot02, Théorème 8], which relies on a study of the properties of the set of zeroes of $Q_{1}^{2}Q_{2}-Q_{1}Q_{2}^{2}$, the approach of Proposition 4.20 is of a significantly different nature.
For $n>2$, the approach of Remark 4.21 is not enough to give a complete classification, but we can extract sufficient conditions for the existence of spaces $L_{2^{n-1}\lambda,n}$ that lead to the discovery of new large classes of examples for every $n$.
Proposition 4.22.
Let $p=2$ and $n\geq 3$. Let $\underline{Q}\in k[X]^{n}$ be a $n$-tuple giving rise to a basis of a space $\Omega$ as in Definition 4.4. For every $i=1,\dots,n$ let $U_{i},V_{i}\in k[X]$ be such that $Q_{i}=U_{i}^{2}+XV_{i}^{2}$ and assume that they satisfy the system of equations